# The Honest Truth About Causal Trees: Accuracy Limits for Heterogeneous Treatment Effect Estimation

### v1

@@ Abstract
Recursive decision trees have emerged as a leading methodology for heterogeneous causal treatment effect estimation and inference in experimental and observational settings. These procedures are fitted using the celebrated CART (Classification And Regression Tree) algorithm (Breiman et al., 1984), or custom variants thereof, and hence are believed to be “adaptive” to high-dimensional data, sparsity, or other specific features of the underlying data generating process. Athey and Imbens (2016) proposed several “honest” causal decision tree estimators, which have become the standard in both academia and industry. We study their estimators, and variants thereof, and establish lower bounds on their estimation error. We demonstrate that these popular heterogeneous treatment effect estimators cannot achieve a polynomial-in-$n$ convergence rate under basic conditions, where $n$ denotes the sample size. Contrary to common belief, honesty does not resolve these limitations and at best delivers negligible logarithmic improvements in sample size or dimension. As a result, these commonly used estimators can exhibit poor performance in practice, and even be inconsistent in some settings. Our theoretical insights are empirically validated through simulations.
Keywords: recursive partitioning, decision trees, causal inference, heterogeneous treatment effects
@@ 1 Introduction
Athey and Imbens (2016) proposed to use recursive decision trees to estimate (and later conduct inference about) heterogeneous causal effects in experimental and observational settings. Their methodology is often called “honest” causal trees. Due in part to its simple, interpretable structure, their causal inference methodology has been widely adopted in academic and industry empirical research over the last decade. For example, to advocate for their proposal, the authors wrote that “[i]t enables researchers to let the data discover relevant subgroups while preserving the validity of confidence intervals constructed on treatment effects within subgroups” (Athey and Imbens, 2016, page 7353).
Despite the widespread use of honest causal tree estimators, little is known about their theoretical properties for estimation and inference. Existing results typically require very strong assumptions on the tree-growing process (Wager and Athey, 2018), which we show are incompatible with canonical implementations of causal trees under basic conditions. Specifically, this paper establishes lower bounds on the estimation error of heterogeneous treatment effect estimators based on recursive adaptive partitioning. We demonstrate that such estimators cannot achieve a polynomial-in-$n$ convergence rate under basic conditions, where $n$ denotes the sample size. Instead, these popular estimators can exhibit arbitrarily slow convergence rates, if not become inconsistent in some cases. As a consequence, our theoretical insights demonstrate that honest causal tree estimators, and variant thereof, may be inaccurate for estimating heterogeneous causal effects, and invalid for constructing confidence intervals on treatment effects within subgroups.
@@
Our work in the causal setting also complements the rich existing theoretical analyses of recursive adaptive partitioning estimators for regression (Scornet et al., 2015; Chi et al., 2022; Klusowski and Tian, 2024; Cattaneo et al., 2024; Mazumder and Wang, 2024) and contributes to the small but growing body of negative results. For example, Ishwaran (2015) showed that regression trees via CART methodology (Breiman et al., 1984) can create imbalanced cells containing a small number of samples. Tan et al. (2022) proved that regression trees are inefficient at estimating additive structure, regardless of the way in which they are optimized. Tan et al. (2024b) proved that mixing times for Bayesian Additive Regression Trees (BART) (Chipman et al., 2010) can increase with the training sample size. Finally, Tan et al. (2024a) established that adaptive regression trees with Boolean covariates can require exponentially many samples in the dimension and are high-dimensional inconsistent for learning ANOVA decompositions with certain interaction patterns.
The present paper supersedes the unpublished manuscript by Cattaneo, Klusowski, and Tian (2022), which showed that a one-dimensional regression stump (i.e., single-split regression trees with a single covariate) constructed via CART can suffer arbitrarily slow convergence rates, and furthermore conjectured (but did not prove) that causal trees might (i) exhibit the same pathology and (ii) fail to benefit from honesty. Our paper proves both conjectures, and goes further by establishing these results for arbitrary covariate dimension and for any causal tree structure with at least one split (i.e., allowing for an arbitrary number of splits or depth of the causal tree). The supplemental appendix also reports analogous results for plain adaptive regression trees. As sketched in Section 4.1, with full details given in the supplemental appendix, our method of proof relies on new insights concerning non-asymptotic approximations for the suprema of partial sums and various Gaussian processes, which may be of independent theoretical interest. In particular, we correct an error in Eicker (1979).
@@ 2 Setup
The available data $\mathcal{D}=\{(y_{i},\mathbf{x}^{\top}_{i},d_{i}):i=1,2,\dots,n\}$ is a random sample, where $y_{i}$ is an outcome variable, $\mathbf{x}_{i}=(x_{1,1},\ldots,x_{1,p})^{\top}$ is a vector of (pre-treatment) covariates, and $d_{i}$ is a binary treatment indicator. Employing standard potential outcomes notation (see, e.g., Hernán and Robins, 2020, for an introduction), we assume that
@@
$\tau(\mathbf{x})\equiv\mathbb{E}\big{[}y_{i}(1)-y_{i}(0)\big{|}\mathbf{x}_{i}=\mathbf{x}\big{]},$
which captures average treatment effects for different values of observable (pre-treatment) covariates. In experimental settings, the CATE function is identifiable because
$\tau(\mathbf{x})$ $=\mathbb{E}\big{[}y_{i}\big{|}d_{i}=1,\;\mathbf{x}_{i}=\mathbf{x}\big{]}-\mathbb{E}\big{[}y_{i}\big{|}d_{i}=0,\;\mathbf{x}_{i}=\mathbf{x}\big{]}$ (1)
@@
where the probability of treatment assignment $\xi=\mathbb{P}(d_{i}=1)$ is known by virtue of the known randomization mechanism. The first equality (1) represents $\tau(\mathbf{x})$ as the difference of two conditional expectation functions based on observed data, while the second equality (2) represents $\tau(\mathbf{x})$ as a single conditional expectation of the “transformed” outcome $y_{i}\frac{d_{i}-\xi}{\xi(1-\xi)}$.
Traditional semiparametric methods would replace the unknown conditional expectations by estimators thereof to learn about heterogeneous treatment effects from experimental data. These methods do not cope well with high-dimensional data, sparsity, or other unknown specific features of the data generating process. Motivated by the recent success of modern (adaptive) machine learning methods, Athey and Imbens (2016) proposed to estimate $\tau(\mathbf{x})$ using recursive decision trees. While retaining the core ideas underlying the greedy recursive construction via standard CART, their proposals customized the tree splitting criterion to the causal inference setting, and employed sample splitting (the so-called “honesty” property) to de-couple the tree construction from the estimation of $\tau(\mathbf{x})$ on the terminal nodes of the tree. This honesty modification has been viewed as a natural “fix,” since separating model selection from estimation is believed to reduce overfitting and improve the validity of inference. Despite this prevailing view, we show that honesty cannot overcome the fundamental limitations of recursive partitioning for heterogeneous causal effect estimation (or for plain adaptive regression trees), offering only at best negligible logarithmic improvements in sample size or dimension.
We perform a comprehensive study of the estimation accuracy of nine distinct causal tree methods, which differ on how their three key underlying parts are implemented: (i) CATE estimator, (ii) tree construction, and (iii) sample splitting.
@@ 2.2 Tree Construction
The variance maximization splitting criterion is somewhat different than the original CART criteria (3), since it explicitly selects splits based on maximizing the squared difference of the child treatment effect estimates. For the $\mathtt{IPW}$ estimator, this rule is equivalent to applying the CART criterion in (3) to the transformed outcome $\tilde{y}_{i}=y_{i}\dfrac{d_{i}-\xi}{\xi(1-\xi)}$. This transformation satisfies $\mathbb{E}[\tilde{y}_{i}\mid\mathbf{x}_{i}=\mathbf{x}]=\tau(\mathbf{x})$ for all $\mathbf{x}\in\mathcal{X}$, and thus CART operates on an outcome whose conditional mean equals the CATE. The $\mathtt{DIM}$ estimator follows the same idea of predicting the within-node average treatment effect, but it constructs these predictions somewhat differently.
@@
Each of the causal recursive tree constructions leads to a distinct data-driven partition of $\mathcal{X}$. A key observation in this paper is that they do not generate quasi-uniform partitions, and thus known results in the nonparametric partitioning-based estimation literature (Györfi et al., 2002; Cattaneo et al., 2020, 2025) are not applicable. The supplemental appendix considers other recursive partitioning constructions, including the standard CART algorithm and variants thereof.
@@ 2.3 Sample Splitting
The final ingredient of the causal tree estimators concerns the data used at each stage of their construction. It is believed that de-coupling the CATE estimation (Definition 1) from the tree implementation (Definition 2) can lead to better performance of the final estimator. In practice, this approach corresponds to sample splitting, and Athey and Imbens (2016) and others referred to it as “honesty.” To avoid confusion, we emphasize that procedures without sample splitting are not “dishonest” in any formal sense; they are simply harder to analyze formally.
@@ 3 Assumptions
We impose the following assumption throughout the paper.
@@
Assumption 1(i) corresponds to simple randomized experiments. Assumption 1(ii) further assumes a canonical homoskedastic causal regression model, while Assumption 1(iii) implies that there is no heterogeneity in the causal treatment effect $\tau=c_{1}-c_{0}$. Because trees are invariant with respect to monotone transformations of the coordinates of $\mathbf{x}_{i}$, without loss of generality, Assumption 1(iv) can be replaced by the assumption that covariates are uniformly distributed on $\mathcal{X}=[0,1]^{p}$, i.e., $x_{i,j}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\text{Uniform}([0,1])$ for $j=1,2,\dots,p$. Finally, Assumption 1(v) means that potential outcome errors are sub-exponential, or equivalently, they satisfy a Bernstein moment condition.
Since we are interested in establishing lower bounds on the estimation accuracy of the causal tree estimators in Definition 3, it is sufficient to consider the constant treatment effect model in Assumption 1 for several reasons. First, this statistical model is a canonical member of any interesting class of data generating processes because the constant function belongs to all classical smoothness function classes, as well as to the set of functions with bounded total variation. It follows that our results will shed light in settings where uniformity over any of the aforementioned classes of functions is of interest: our lower bounds can be applied directly in those cases because for any estimator $\hat{\tau}(\mathbf{x})$ of the parameter $\tau(\mathbf{x})$,
@@
Second, Assumption 1 also removes issues related to smoothing (or misspecification) bias, heteroskedasticity, and heavy tail distributions. In particular, since the CATE function $\tau(\mathbf{x})$ is constant for all $\mathbf{x}\in\mathcal{X}$, our results will not be driven by standard (boundary or other smoothing) bias in nonparametrics. For example, if the distributions of $\varepsilon_{i}(0)$ and $\varepsilon_{i}(1)$ are symmetric about zero,
@@
for $l\in\{\mathtt{DIM},\mathtt{IPW},\mathtt{SSE}\}$ and $\mathbf{x}\in\mathsf{t}$ where $\mathsf{t}$ is a terminal node in the tree. Unbiasedness of $\hat{\tau}^{\mathtt{NSS}}_{l}(\mathbf{x})$ follows from the fact that the split points are symmetric functions of the residuals. In the case of $\hat{\tau}^{\mathtt{HON}}_{l}(\mathbf{x})$, sample splitting can generate empty cells with positive probability, which is captured by the term $\tau\mathbb{P}(n(\mathsf{t})=0)$; see Lemma SA-37 in the supplemental appendix. It follows that, in particular, $\hat{\tau}^{\mathtt{HON}}_{l}(\mathbf{x})$ is unbiased when $\tau=0$ (or for any other known treatment effect value), as well as in tree constructions ensuring that $\mathbb{P}(n(\mathsf{t})=0)=0$; otherwise, $\hat{\tau}^{\mathtt{HON}}_{l}(\mathbf{x})$ is asymptotically unbiased whenever $\mathbb{P}(n(\mathsf{t})=0)\to 0$ as $n\to\infty$. Our results will be driven by the fact that canonical adaptive decision tree constructions can generate small cells containing only a handful of observations, thereby making the estimator highly inaccurate in some regions of $\mathcal{X}$, regardless of bias. In other words, inconsistency is due to a large variance problem, not a large bias problem.
Third, the local constant treatment effect model could also be interpreted as a first-order approximation of the smooth function $\tau(\mathbf{x})$. Because the recursive partitioning schemes lead to a partitioning-based estimator of the CATE function, it follows that $\tau(\mathbf{x})$ is approximated locally by a Haar basis (piecewise constant functions). In fact, our results can be extended to hold uniformly over appropriate shrinking neighborhoods of smooth functions local to the constant function, provided that the signal to noise ratio (bias-variance trade-off) is small.

### v2

@@ Abstract
Recursive decision trees are widely used to estimate heterogeneous causal treatment effects in experimental and observational studies. These methods are typically implemented using CART-type recursive partitioning and are often viewed as adaptive procedures capable of discovering treatment effect heterogeneity in high-dimensional settings. We study causal tree estimators based on adaptive recursive partitioning and establish lower bounds on their estimation accuracy. Under basic conditions, we show that causal trees constructed via standard CART-type splitting rules cannot achieve polynomial-in-$n$ convergence rates in the uniform norm (where $n$ denotes the sample size). The underlying mechanism is that greedy recursive partitioning selects highly imbalanced splits with non-vanishing probability, producing terminal nodes containing very few observations and leading to large estimation variance. We further show that sample splitting (“honesty”) yields at most negligible improvements in convergence rates. As a consequence, causal tree estimators may converge arbitrarily slowly and can even be inconsistent in some settings. Our results also clarify the role of balanced partition assumptions in existing theoretical guarantees for causal forests and related ensemble methods. The analysis develops new probabilistic tools for studying adaptive recursive partitioning procedures, including non-asymptotic approximations for suprema of partial sums and Gaussian processes. As a technical by-product, we also identify and correct an error in Eicker (1979).
Keywords: machine learning, recursive partitioning, adaptive decision trees, minimax convergence rates
Contents
@@ 1 Introduction
Recursive decision trees have become a popular tool for estimating heterogeneous causal effects in both experimental and observational settings. These methods adapt the classical CART (Classification and Regression Tree) algorithm (Breiman et al., 1984) to causal inference by modifying the splitting criterion and (sometimes) using sample splitting to separate tree construction from treatment effect estimation. Due to their simplicity and interpretability, causal trees have been widely adopted in both academic research and industry applications over the last decade. A leading example is the honest causal tree framework of Athey and Imbens (2016), which is often understood as allowing the data to “discover relevant subgroups while preserving the validity of confidence intervals constructed on treatment effects within subgroups” (Athey and Imbens, 2016, page 7353).
Despite their widespread use, the fundamental statistical properties of causal tree estimators and their associated inference procedures remain poorly understood. Existing theoretical analyses typically rely on strong assumptions about the tree-growing process that are incompatible with canonical implementations of CART-type recursive partitioning under basic conditions. This paper studies heterogeneous treatment effect estimators based on adaptive recursive partitioning and establishes theoretical lower bounds on their estimation accuracy. More broadly, our results reveal a fundamental limitation of CART-type recursive partitioning for uniform estimation of heterogeneous treatment effects.
Our main finding is that causal tree estimators constructed via standard CART-type greedy splitting cannot achieve polynomial-in-$n$ convergence rates in the uniform norm under basic conditions, where $n$ denotes the sample size. More precisely, we show that these estimators converge more slowly than any polynomial-in-$n$ rate uniformly over the covariate space. This failure arises even in the simplest setting in which the true treatment effect is constant over the feature space.
The mechanism behind this phenomenon is intrinsic to greedy recursive partitioning. CART-type splitting rules select highly imbalanced splits with non-vanishing probability, thereby generating terminal nodes with very few observations. Because treatment effects within each node are estimated by local averaging, these small cells induce large estimation variance and prevent the estimator from concentrating uniformly around the true conditional average treatment effect.
A natural response to this small-cell phenomenon is to impose additional regularization on the tree construction. In practice, this often takes the form of minimum node-size requirements. Although such modifications may reduce finite-sample variance, they also alter the class of feasible partitions of the procedure. In particular, ruling out small cells necessarily introduces additional approximation bias whenever the conditional average treatment effect (CATE) is not locally constant, and this problem can be even more pronounced under conditional heteroskedasticity. More broadly, the more aggressively one regularizes the tree through minimum node-size requirements, the more one weakens the adaptivity that makes CART attractive in the first place. For this reason, in practice, causal tree methods are typically regularized through fixed or lightly tuned minimum node-size hyperparameters, often on the order of only a few tens of observations, rather than through carefully tuned sample-size-dependent hyperparameters that grow with $n$, as would be required to address the kind of convergence issues raised here.
At present, however, the statistical behavior of CART-type estimators under such algorithmic regularization remains only partially understood in general settings where the partition is selected adaptively from the data. For this reason, regularization alone does not provide a theoretically satisfactory resolution of the phenomenon we document unless one can formally characterize the resulting bias–variance trade-off and establish valid inference for the resulting estimator. Moreover, even apart from these statistical concerns, it has been argued that imbalanced splits can be algorithmically beneficial for deeper trees, because they preserve sample size after a bad split and allow the tree to recover downstream (Ishwaran, 2015).
Another frequently discussed modification of adaptive recursive partitioning methods is sample splitting, also known as honesty. By constructing the tree on one subsample and estimating treatment effects on an independent subsample, honesty aims to reduce overfitting and improve the validity of inference. While honesty does yield modest improvements in the achievable convergence rate, our results demonstrate that it does not resolve the fundamental limitations of adaptive recursive partitioning. Even under honest sample splitting, causal tree estimators cannot achieve polynomial-in-$n$ uniform convergence rates under basic conditions.
Our findings also have implications for theoretical guarantees established for causal forests and related ensemble methods. Existing polynomial-rate analyses of forest estimators often rely on conditions ensuring that each constituent tree generates approximately balanced partitions. These conditions are commonly formalized through the so-called $\alpha$-regularity assumption (Wager and Athey, 2018), requiring each split to allocate a non-negligible fraction of the observations in the parent node to each child node. While such assumptions facilitate theoretical analysis, they are incompatible with canonical CART-type greedy splitting rules: under basic conditions, CART constructions can select highly imbalanced splits with non-vanishing probability. Consequently, theoretical guarantees for causal forests that rely on quasi-uniform partitions or $\alpha$-regularity do not directly extend to implementations based on standard CART-type splitting procedures. Moreover, even within the $\alpha$-regular framework itself, these guarantees lead to a counterintuitive conclusion. In the class analyzed by Wager and Athey (2018, see Theorem 3.2 and condition (14) in Theorems 3.1 and 4.1), both the bias exponent and the subsampling-rate exponent improve monotonically with $\alpha$. The available upper bounds therefore favor increasingly balanced splits. If one extrapolates these expressions over $\alpha\in(0,1/2]$, the most favorable bound is achieved at $\alpha=1/2$, that is, at the maximally balanced case. See Section 6 for further discussion.
From an inferential perspective, the same mechanism underlying our lower bounds also creates a fundamental obstacle for subgroup-level inference based on causal trees and their ensembles. In regions where terminal nodes contain very few observations, the effective sample size may fail to increase with the overall sample size. In such settings, classical Gaussian approximations and standard error formulas need not be valid. Establishing distributional approximations for CART-based causal tree or forest estimators without imposing quasi-uniform partition conditions therefore remains an open theoretical problem.
Before introducing the formal setup, it is useful to briefly summarize the intuition behind our results. Recursive decision trees select splits by optimizing a data-dependent criterion over many candidate partitions of the covariate space. When the underlying conditional expectation function is locally flat, as in the constant treatment effect model considered in this paper, the splitting criterion is driven primarily by stochastic fluctuations in the data. As a consequence, with non-vanishing probability the optimal split occurs near the boundary of a parent node, producing highly imbalanced child nodes. Because treatment effects within each node are estimated by local averaging, such splits generate terminal nodes containing very few observations, leading to large estimation variance in those regions of the covariate space. The recursive nature of the tree construction then propagates this phenomenon across deeper levels of the tree, creating multiple regions with small effective sample sizes and ultimately preventing uniform convergence of the estimator.
1.1 Contributions and Related Literature
This paper makes three main contributions. First, we establish lower bounds on the uniform convergence rate of causal tree estimators constructed via CART-type recursive partitioning. These results show that such estimators cannot achieve polynomial-in-$n$ convergence rates under basic conditions, even in settings where the underlying treatment effect is constant. Second, we analyze the role of sample splitting (honesty) in causal tree estimation. While honesty improves the achievable convergence rate by removing a slowly varying $\sqrt{\log\log n}$ factor, our results demonstrate that it does not eliminate the fundamental limitations of adaptive recursive partitioning. Third, we clarify the implications of these findings for tree-based causal inference more broadly, including commonly used regularization strategies and theoretical guarantees for causal forests that rely on quasi-uniform partitions or $\alpha$-regularity conditions. Our analysis shows that these assumptions are incompatible with canonical CART-type implementations and highlights the resulting challenges for valid inference based on tree-grown partitions.
@@ 1.1 Contributions and Related Literature
Our work complements the rich theoretical literature on recursive adaptive partitioning estimators for regression (Scornet et al., 2015; Chi et al., 2022; Klusowski and Tian, 2024; Cattaneo et al., 2024; Mazumder and Wang, 2024) and contributes to a growing body of negative results for tree-based methods. For example, Tan et al. (2022) demonstrated that regression trees are inefficient at estimating additive structure, regardless of the optimization strategy employed. Tan et al. (2024) established that the mixing time of Bayesian Additive Regression Trees (BART) (Chipman et al., 2010) can increase with the training sample size. Finally, Tan et al. (2026) showed that adaptive regression trees with Boolean covariates may require exponentially many samples in the dimension and can be inconsistent in high-dimensional settings.
The present paper supersedes the unpublished manuscript of Cattaneo, Klusowski, and Tian (2022), which studied the behavior of a one-dimensional regression stump constructed via CART and showed that its estimation error can converge arbitrarily slowly. That work conjectured that similar phenomena would arise for causal trees and that honesty would not eliminate the problem. The present paper confirms both conjectures and substantially generalizes those findings. In particular, we establish lower bounds for causal tree estimators in arbitrary covariate dimension and for any causal tree structure containing at least one split, allowing for arbitrary tree depth. Furthermore, the supplemental appendix reports analogous results for standard adaptive regression trees of arbitrary depth, and with and without sample splitting. These theoretical results clarify the role of necessary and sufficient conditions in the work of Bühlmann and Yu (2002) and Banerjee and McKeague (2007), among others. See Section 6 for more discussion.
The proofs rely on a collection of new probabilistic insights concerning the behavior of adaptive recursive partitioning procedures. Our analysis develops non-asymptotic approximations for suprema of partial sums and related Gaussian processes, leveraging high-dimensional central limit theorems, Gaussian comparison inequalities, and strong approximation techniques. Key technical tools include results from Chernozhukov et al. (2017), Chernozhuokov et al. (2022), Csörgö and Révész (1981), Csörgö and Horváth (1997), El-Yaniv and Pechyony (2009), Horváth (1993), Göing-Jaeschke and Yor (2003), Latała and Matlak (2017), Petrov (2007), Shorack and Smythe (1976), and Skorski (2023). As a technical by-product of our analysis, we also identify and correct an error in Eicker (1979).
1.2 Organization
The remainder of the paper is organized as follows. Section 2 introduces the causal tree estimators and the data generating framework studied in the paper. Section 3 presents the assumptions underlying our theoretical analysis. Section 4 establishes lower bounds on the uniform convergence rate of causal tree estimators and provides complementary results on their integrated mean squared error. Section 5 studies $\mathbf{X}$-adaptive variants of causal tree estimators and demonstrates their uniform inconsistency under increasing tree depth. Section 6 discusses the implications of our results for recursive partitioning methods, causal forests, and inference procedures. Section 7 presents simulation evidence illustrating the practical implications of the theoretical results. Appendix A identifies and corrects an error in Eicker (1979). Additional theoretical results and all technical proofs are reported in the supplemental appendix.
@@ 2 Setup
We introduce the class of causal tree estimators studied in the paper. These estimators combine three components: (i) a within-node estimator of the conditional average treatment effect (CATE), (ii) a recursive partitioning rule used to construct the tree, and (iii) a data usage scheme determining whether sample splitting (honesty) is employed. Different choices of these components lead to the family of estimators analyzed in this paper.
The available data $\mathcal{D}=\{(y_{i},d_{i},\mathbf{x}^{\top}_{i}):i=1,2,\dots,n\}$ is a random sample, where $y_{i}$ is an outcome variable, $\mathbf{x}_{i}=(x_{i,1},\ldots,x_{i,p})^{\top}$ is a vector of pre-treatment covariates, and $d_{i}$ is a binary treatment indicator. Employing standard potential outcomes notation (see, e.g., Hernán and Robins, 2020), we assume that
@@
$\tau(\mathbf{x})=\mathbb{E}\big[y_{i}(1)-y_{i}(0)\mid\mathbf{x}_{i}=\mathbf{x}\big],$
which captures how treatment effects vary with observable pre-treatment covariates. In experimental settings, the CATE function is identifiable because
$\tau(\mathbf{x})$ $=\mathbb{E}\big[y_{i}\mid d_{i}=1,\;\mathbf{x}_{i}=\mathbf{x}\big]-\mathbb{E}\big[y_{i}\mid d_{i}=0,\;\mathbf{x}_{i}=\mathbf{x}\big]$ (1)
@@
where the probability of treatment assignment $\xi=\mathbb{P}(d_{i}=1)$ is known by virtue of the known randomization mechanism. The first equality (1) expresses the CATE as the difference of two conditional expectation functions based on observed data, while the second equality (2) represents the CATE as a single conditional expectation of the transformed outcome $y_{i}\frac{d_{i}-\xi}{\xi(1-\xi)}$.
Traditional semiparametric approaches estimate heterogeneous treatment effects by replacing the conditional expectations in (1) or (2) with nonparametric estimators. However, such methods may perform poorly in high-dimensional settings or when the structure of the regression functions (e.g., sparsity or additive separability) is unknown. Motivated by the recent success of adaptive machine learning methods, Athey and Imbens (2016) proposed estimating $\tau(\mathbf{x})$ using recursive decision trees. Their approach retains the greedy recursive construction of standard CART while modifying the splitting criterion to target treatment effect heterogeneity. They also proposed using sample splitting (the so-called “honesty” property) to decouple tree construction from the estimation of treatment effects within terminal nodes.
The introduction of honesty has been widely viewed as a natural “fix,” since separating model selection from estimation is believed to reduce overfitting and improve the validity of inference. Despite this prevailing view, we show that honesty does not overcome the fundamental limitations of recursive partitioning for heterogeneous causal effect estimation (or even for adaptive regression trees), yielding at best negligible logarithmic improvements in the effective sample size or dimension.
We study causal tree estimators obtained by combining three components: two within-node treatment effect estimators (difference-in-means and inverse probability weighting), two tree construction criteria (variance maximization and SSE minimization), and two sample splitting schemes (with and without honesty). These combinations yield nine causal tree estimators corresponding to commonly used implementations in practice.
@@ 2.2 Tree Construction
The variance maximization splitting criterion differs from the original CART criterion (3) in that it explicitly selects splits that maximize the squared difference between the estimated treatment effects in the two child nodes. In this sense, the rule directly targets treatment effect heterogeneity across the partition. For the $\mathtt{IPW}$ estimator, this rule is equivalent to applying the CART criterion in (3) to the transformed outcome $\tilde{y}_{i}=y_{i}\dfrac{d_{i}-\xi}{\xi(1-\xi)}$. This transformation satisfies $\mathbb{E}[\tilde{y}_{i}\mid\mathbf{x}_{i}=\mathbf{x}]=\tau(\mathbf{x})$ for all $\mathbf{x}\in\mathcal{X}$, and thus CART operates on an outcome whose conditional mean equals the CATE. The $\mathtt{DIM}$ estimator follows the same idea of predicting the within-node average treatment effect, but it constructs these predictions somewhat differently.
@@
Each of the causal recursive tree constructions leads to a distinct data-driven partition of $\mathcal{X}$. A key observation underlying our analysis is that these recursive procedures do not generate quasi-uniform partitions of the covariate space, and thus known results in the nonparametric partitioning-based estimation literature (Györfi et al., 2002; Cattaneo et al., 2020, 2026) are not applicable. The supplemental appendix considers other recursive partitioning constructions, including the standard CART algorithm and variants thereof.
@@ 2.3 Sample Splitting
The final ingredient of the causal tree estimators concerns the data used at each stage of their construction. It is believed that de-coupling the CATE estimation (Definition 1) from the tree implementation (Definition 2) can lead to better performance of the final estimator. In practice, this approach corresponds to sample splitting, and Athey and Imbens (2016) and others referred to it as “honesty.” To avoid confusion, procedures without sample splitting are not “dishonest” in any formal sense; they are simply harder to analyze theoretically.
@@
Recall Definition 1 and Definition 2, and that $\mathcal{D}=\{(y_{i},d_{i},\mathbf{x}^{\top}_{i}):i=1,2,\dots,n\}$ is the available random sample.
@@ 3 Assumptions
The following assumption describes the data generating process used throughout the analysis.
@@
Assumption 1(i) corresponds to a simple randomized experiment with treatment probability $\xi\in(0,1)$. Assumption 1(ii) specifies a canonical regression representation for the potential outcomes, where $\mu_{d}(\mathbf{x})$ denotes the conditional mean function and $\varepsilon_{i}(d)$ is a mean-zero error term independent of $\mathbf{x}_{i}$. Assumption 1(iii) imposes a constant treatment effect, $\tau=c_{1}-c_{0}$, across the covariate space. Assumption 1(iv) requires the covariates to be independent and continuously distributed. Because recursive decision trees are invariant to monotone transformations of the covariates, this condition can be replaced without loss of generality by assuming that $\mathbf{x}_{i}$ is uniformly distributed on $\mathcal{X}=[0,1]^{p}$. Finally, Assumption 1(v) requires the potential outcome errors to be sub-exponential, or equivalently to satisfy a Bernstein moment condition.
Because our goal is to establish lower bounds on the estimation accuracy of the causal tree estimators defined in Definition 3, it suffices to consider the constant treatment effect model in Assumption 1. The constant function belongs to all commonly studied smoothness classes, including Hölder classes and functions with bounded total variation. Consequently, lower bounds established under this model immediately extend to larger classes of data generating processes. Formally, for any estimator $\hat{\tau}(\mathbf{x})$ and any class of distributions $\mathcal{P}$ containing the distribution $\mathbb{P}_{1}$ satisfying Assumption 1,
@@
Assumption 1 also removes issues related to smoothing (or misspecification) bias, heteroskedasticity, and heavy-tailed errors. In particular, because the CATE function $\tau(\mathbf{x})$ is constant over $\mathcal{X}$, the results below are not driven by the usual (boundary or smoothing) bias arising in nonparametric estimation. For example, if the distributions of $\varepsilon_{i}(0)$ and $\varepsilon_{i}(1)$ are symmetric about zero, then
@@
for $l\in\{\mathtt{DIM},\mathtt{IPW},\mathtt{SSE}\}$ and $\mathbf{x}\in\mathsf{t}$ where $\mathsf{t}$ is a terminal node in the tree. Unbiasedness of $\hat{\tau}^{\mathtt{NSS}}_{l}(\mathbf{x})$ follows from the fact that the split points are symmetric functions of the residuals. In the case of $\hat{\tau}^{\mathtt{HON}}_{l}(\mathbf{x})$, sample splitting can generate empty cells with positive probability, which is captured by the term $\tau\mathbb{P}(n(\mathsf{t})=0)$; see Lemma SA-37 in the supplemental appendix. In particular, $\hat{\tau}^{\mathtt{HON}}_{l}(\mathbf{x})$ is unbiased when $\tau=0$ (or for any other known treatment effect value), as well as in tree constructions ensuring that $\mathbb{P}(n(\mathsf{t})=0)=0$. More generally, $\hat{\tau}^{\mathtt{HON}}_{l}(\mathbf{x})$ is asymptotically unbiased whenever $\mathbb{P}(n(\mathsf{t})=0)\to 0$ as $n\to\infty$.
Consequently, the negative results established below are not driven by bias. Instead, they arise because adaptive recursive tree constructions can generate highly imbalanced partitions with non-vanishing probability, producing terminal nodes that contain very few observations. These small cells lead to large estimation variance in some regions of $\mathcal{X}$, which ultimately prevents the estimator from achieving polynomial convergence rates.
Finally, the constant treatment effect model can also be interpreted as a local approximation to smooth heterogeneous treatment effect functions. Recursive partitioning estimators approximate $\tau(\mathbf{x})$ using piecewise constant functions over the tree partition, which corresponds to a Haar basis representation. Our results therefore extend to shrinking neighborhoods of smooth functions around the constant function when the signal-to-noise ratio is sufficiently small.